OpenStudy (anonymous):

will give medal !!!!!!!! Becca's family has 4 gallons of milk. They drink 5/8 gallon per day. How long will their milk last? A. 2 1/2 days B. 4 days C. 4 5/8 days D. 6 2/5 days

OpenStudy (anonymous):

@Ace8407 it's more helpful if you explain how you got the answer :)

hero (hero):

It's really sad that users like @Ace8407 are willing to give an answer that is against CoC but refuse to explain anything When the explanation and guidance is the kind of help the asker really needs.

OpenStudy (anonymous):

To find the number of days (n) we divide the total amount of milk by the amount consumed per day. n = 4 ÷ 8/5 When dividing by a fraction, this is the same as multiplying by the upside down version of the fraction. So, 4 ÷ 5/8 = 4 x 8/5 Multiply the numerators, and; =32/5 Simplify; 6 2/5

OpenStudy (anonymous):

Then he should have said "can you explain your answer" so stfu. and how do u know I wasn't going to message him and tell him how I got it just because I realized that I probably should after I said " no thanks" to @livias.random so don't talk. @Hero

OpenStudy (anonymous):

Give Help, Not Answers: I will encourage and guide those needing help, and not just give them an answer (Don’t post only answers - guide the asker to a solution.) - From the Code of Conduct

hero (hero):

@Ace8407 1. Read the Code of Conduct: http://openstudy.com/code-of-conduct 2. Acronyms like the kind you just used are considered profanity. 3, On OS we appeaate your help, but as a helper you are expected to give hints, explanations, and guidance rather than just answers. 4. It is best if you can help the asker figure out how to solve the problem on their own.

OpenStudy (anonymous):

k